# From User Surveys to Telemetry-Driven Agents: Exploring the Potential of Personalized Productivity Solutions

### v1

@@ 3.1. Understanding User Preferences for Productivity Agents
The primary objective of our first survey was to identify the ideal persona for a productivity agent by examining people’s expectations and preferences concerning its features and behaviors. We conducted a user survey with N=363 information workers, using this as an initial step to gather insights into their desired characteristics and capabilities for a productivity agent. The survey delved into various design aspects, including the agent’s empathy level, proactive assistance, communication style, and adaptability to different situations. The survey we designed was carefully crafted based on prior work  (Grover et al., 2020; Kimani et al., 2019) and extensive interactions with experts in the field (i.e., AI professionals, organizational psychologists with expertise in workplace productivity and technology adoption, productivity engineers and expert HCI researchers). Drawing upon their knowledge and experience, we aimed to create a comprehensive and insightful questionnaire that would effectively capture users’ preferences and concerns in relation to a productivity agent. Refer to Appendix A for a comprehensive list of questions posed to participants.
@@ 3.2. Productivity Solutions Design and Evaluation
To personalize the experience to each individual, we decided to use the high-level telemetry data on participants’ meetings, communication, collaboration network, and other work habits which are provided through Microsoft’s Viva Insights [^1] https://learn.microsoft.com/en-us/viva/insights/Use/Metric-definitions. Among the different platforms, we selected Viva Insights as it is widely used in many companies and computes a wide variety of relevant features in the context of productivity. To extract the information, we built a custom dashboard scraper (see Figure 2) that allowed participants to automatically extract their Viva Insights data, review and modify it, and ultimately share it with the researchers. The extracted data was then used to create the three modes of presentation tailored to each participant’s work patterns and habits. After participants willingly shared their telemetry data with us, they engaged with the three modes: viewing their data on a dashboard, reading a personalized narrative, and interacting with AI productivity agent that had access to their telemetry data. They then assessed their experiences based on the ability to gain productivity insights or identify opportunities for productivity improvement by answering our follow-up survey. Note, the AI productivity agent primarily functioned as a ‘productivity guide,’ offering recommendations and advice to improve users’ productivity. This included suggesting strategies for time management, prioritizing tasks, and optimizing work routines. However, it is important to note that the agent did not directly perform tasks such as meeting summarization or documentation writing. While it provided guidance in these areas, the execution of tasks remained the user’s responsibility. This distinction is crucial for understanding the agent’s role in our study. In total, we had N=40 participants complete the second phase of our study. To diversify our sample, we reached out to individuals who had not participated in the first phase of the study. The average time for study completion was 30.15 minutes, and participants received a \$25 gift card as compensation. Our study exclusively makes use of Open AI’s  (OpenAI, 2023b) GPT-4 API.
@@
Telemetry Data Collection: Our scraper tool employs Selenium (Selenium, 2023), a testing library, to open up a web browser and navigate the Viva Insights pages, capturing the HTML source of each page. Since these pages are dynamically generated, we opted for a more error-resilient approach by utilizing GPT-4 to extract the desired information, rather than relying on traditional HTML tag-based scraping. Once we obtained the HTML source, we removed unnecessary HTML tags and attributes and then sent the streamlined source to GPT-4 with a specific prompt to help identify the information we sought. This method proved to be effective in facilitating the extraction of relevant data from the Viva Insights pages. Please refer to Appendix C and Appendix D for the comprehensive list of system prompts and user prompts we provided to GPT-4 for extracting and transforming the desired information. The telemetry information obtained from Viva Insights can be found in Table 1. It is important to note that Viva Insights calculates these insights based on the data from the prior month.
@@
Designing Productivity Solutions: As mentioned earlier, we designed three productivity solutions: an AI productivity agent (also referred to as an interactive chat agent), a personalized narrative (generated via GPT-4), and a traditional visual dashboard. We began by integrating telemetry data directly into the AI agent’s system prompt, a crucial step that empowered the agent with the ability to analyze and respond based on individual user behaviors and patterns. This data-driven approach ensured that each participant received feedback and suggestions that reflected their unique work habits. To ensure the agent’s feedback and recommendations resonated with users, we drew from the insights gathered during the first phase of our study. The preferences and expectations participants shared during this phase became the foundation for designing the system prompt of the AI agent. Parallel to the development of the AI agent, we embarked on creating an AI-based personalized narrative, an alternative means of communicating productivity insights. This narrative was also crafted to align with the specific preferences participants had voiced. Beyond just catering to these preferences, we introduced a structured format to this narrative, ensuring as consistent delivery of insights as we could across participants. The narrative format presented information in a structured, story-like manner, providing users with an overview of their productivity patterns and suggesting improvements. In contrast, the interactive agent engaged users in a conversational interface, offering personalized guidance based on user inputs and queries. This interactive nature allowed for a more tailored and dynamic exchange, though it was limited to offering advice rather than executing tasks directly. To perfect the prompts for both the AI productivity agent and the personalized narrative, we engaged in multiple cycles of testing and refinement. Each iteration was an opportunity to align more closely with user needs and preferences. Refer to Appendix  E for the full prompts used for the productivity agent and personalized narrative. For the development of the dashboard visualization, we recognized that many participants had established familiarity with Viva Insights, so we aimed to make our dashboard’s visual presentations echo those of Viva Insights. We meticulously designed a range of visual representations, from bar charts that provided at-a-glance insights to pie charts that broke down data into comprehensible segments. Furthermore, we chose to present certain telemetry data numerically, ensuring users had access to precise figures. Figure 3 presents the AI productivity agent along with the other two presentation modes.
@@ 4.1. Exploring User Preferences for Productivity Agents
Most of the participants, 169 (60.4%), expressed their preference for the agent to provide proactive suggestions to enhance productivity. In contrast, 77 participants (27.5%) wanted the agent to provide assistance and suggestions only when explicitly asked for help. The agent’s preferred personality traits are depicted in Figure 4(a); the most common preferences were: knowledgeable (88.9%), trustworthy (75.0%), and transparent (68.2%). Similarly, Figure 4(b) represents the social role that people desired for the agent. The top choices were that of an assistant (76.4%) and expert (56.1%). The supervisor role was the least preferred, with only 4.6% of participants opting for it. These findings emphasize the importance of designing a productivity agent with a proactive approach and equipping it with the desired personality traits, such as knowledgeability, trustworthiness, and transparency, to better cater to users’ preferences.
@@
A significant majority of participants, 209 (74.6%), expressed that they would prefer a moderate to high level of personalization in the productivity agent. This enables the agent to provide highly personalized advice tailored to their specific needs, considering various aspects of their work habits and preferences. In terms of privacy, 152 participants (54.3%) felt either somewhat comfortable or completely comfortable sharing their personal information with the productivity agent. However, 58 participants (20.7%) expressed discomfort in sharing their personal information. Interestingly, a majority of participants, 205 (73.2%), were either somewhat comfortable or completely comfortable sharing high-level telemetry data for enhancing the quality and guidance provided by the productivity agent. These findings highlight the importance of striking a balance between personalization and privacy to cater to users’ preferences and concerns.
@@
Productivity Perspectives
Participants felt productive when making progress on scheduled tasks, solving unexpected tasks, and helping colleagues
@@
Agent designed to offer strategies for task management, prioritizing, and collaborative assistance. It is developed to help users achieve an optimal work-life balance and improve their productivity.
Communication Style
Preference for a tone that balances casual and formal styles; adaptability in tone based on task and situation; direct and straightforward communication rather than chatty
@@
Agent uses a professional yet friendly tone, adaptable to the context and user’s current task. It is designed to communicate concisely without being overly chatty or verbose.
Approach
Desire for proactive suggestions and assistance to enhance productivity
@@
Agent designed to provide proactive productivity suggestions without being too intrusive and pushy
Personality Traits
Valued traits included being knowledgeable, trustworthy, and transparent.
@@
Agent developed to demonstrate these traits consistently across interactions.
Personalization
Preference for a moderate to high level of personalization, balancing with privacy concern
@@
Agent uses high-level telemetry data to personalize suggestions while ensuring user privacy
Negative Traits to Avoid
@@
Avoidance of traits like being nagging, aggressive, pushy, overly casual, judgmental, or patronizing.
Agent prompted to avoid these negative traits, focusing on supportive and informative interactions.
@@
4.2. Design of the AI Productivity Agent Informed by User Preferences
We now present how the design of the AI productivity agent was crafted based on the insights obtained from the initial user survey conducted in Phase 1 (Section 4.1). This survey comprehensively explored users’ preferences and expectations in several key areas: productivity perspectives, communication style, approach, personality traits, personalization, and desired and undesired traits in an AI agent. A summary of these findings and their corresponding implementations in the AI agent’s design is presented in Table 2.
The table illustrates how each identified theme from the survey was translated into specific features of the AI productivity agent. For instance, user preferences for a balanced communication style influenced the agent’s development to ensure it communicates in a professional yet approachable manner. Similarly, the agent’s approach and personality traits were tailored to align with users’ desires for proactive assistance and valued characteristics like knowledgeability and trustworthiness. Additionally, personalization emerged as a crucial aspect, with users expressing a need for tailored recommendations balanced against privacy concerns. Accordingly, the agent leverages high-level telemetry data to offer personalized suggestions, carefully designed to respect user privacy. This careful alignment of the AI productivity agent’s features with user preferences underscores our commitment to developing a user-centric tool. It highlights the importance of basing design decisions on thorough user research, ensuring that the developed AI agent is not only technologically advanced but also resonates with the actual needs and preferences of its intended users.
@@ 4.2. Design of the AI Productivity Agent Informed by User Preferences
Figure 6. Overall Experience: Figure (a) illustrates the participants’ overall experience ratings for three data presentation methods: Personalized Narrative, AI Productivity Agent, and Dashboard. Figure (b) compares the ratings of these methods across various criteria. We performed the Kruskal Wallis test to analyze the differences between the three groups in Figure (b), and criteria without any statistically significant differences (i.e., p-value > 0.10) has been marked with a crosshatch shading pattern - a series of intersecting lines drawn at an angle to create texture.
@@
Participants spent an average of 60.58 seconds (with a standard deviation of 38.98 seconds) interacting with the dashboard. In comparison, they spent considerably more time on the AI productivity agent, with an average duration of 315.48 seconds (standard deviation: 272.23 seconds). The personalized narrative had a similar interaction time to the dashboard, with an average of 65.21 seconds (standard deviation: 55.39 seconds). After interacting with each method, participants were asked to rate their overall experience on a six-point Likert scale ranging from Very Poor to Excellent. As shown in the stacked bar graph in Figure 5(a), 26.2% of participants rated the Dashboard as excellent, while only 11.9% and 9.5% rated Excellent for the productivity agent and personalized narrative, respectively. Combining the top three experiences (excellent, very good, and good), 81% considered the Dashboard good or above, 66.7% considered the AI productivity agent good or above, and 64.2% considered the personalized narrative to be good or above. Following the broader question, participants were asked to rate their experience on specific criteria for each method using a five-point Likert scale from Strongly Disagree (1) to Strongly Agree (5). In Figure 5(b), a heatmap displays the average scores obtained by each method for these questions. Most darker blocks (in blue) are centered on the Dashboard, indicating higher average scores for the ease of exploring and understanding data, the use of the presentation method in real-life situations, and enjoyment in interaction. The boxes with crosshatch shading patterns (intersecting lines) indicate that the differences were not statistically significant (Kruskal Wallis test; p-value > 0.10) among the three methods for that particular criterion. The productivity agent scored better in only one area: engagement during data exploration. The personalized narrative performed the worst with the lowest scores among the three. Regarding anticipated usage frequency, we show in Figure 6(a) that 42% of participants selected weekly for dashboard and 30% selected weekly for productivity agent, which was the majority score for both groups. For personalized narrative, 28% participants said that they would use it weekly and the same amount also said they would rather use it monthly. 30% of the participants said they would use the personalized narrative occasionally. Lastly, participants ranked the three methods in order of preference (1 being the most preferred and their first choice, 3 being the least preferred and their last choice). As shown in Figure 6(b), 55% rated the Dashboard as their most preferred method, followed by 27.5% preferring the productivity agent and 17.5% preferring the personalized narrative. Over half of the participants (52.5%) rated the personalized narrative as their least preferred method.

### v2

@@ 3.1. Understanding User Preferences for Productivity Agents
The primary objective of our first survey was to identify the ideal persona for a productivity agent by examining people’s expectations and preferences concerning its features and behaviors. We conducted a user survey with N=363 information workers, using this as an initial step to gather insights into their desired characteristics and capabilities for a productivity agent. The survey focused on various design aspects, including the agent’s empathy level, proactive assistance, communication style, and adaptability to different situations. The survey we designed was carefully crafted based on prior work  (Grover et al., 2020; Kimani et al., 2019) and extensive interactions with experts in the field (i.e., AI professionals, organizational psychologists with expertise in workplace productivity and technology adoption, productivity engineers and expert HCI researchers). Drawing upon their knowledge and experience, we aimed to create a comprehensive and insightful questionnaire that would effectively capture users’ preferences and concerns in relation to a productivity agent. Refer to Appendix A for a comprehensive list of questions posed to participants.
@@ 3.2. Productivity Solutions Design and Evaluation
To personalize the experience for each individual, we decided to use the high-level telemetry data on participants’ meetings, communication, collaboration network, and other work habits which are provided through Microsoft’s Viva Insights [^2] https://learn.microsoft.com/en-us/viva/insights/Use/Metric-definitions. Among the different platforms, we selected Viva Insights as it is widely used in many companies and computes a wide variety of relevant features in the context of productivity. To extract the information, we built a custom dashboard scraper (see Figure 2) that allowed participants to automatically extract their Viva Insights data, review and modify it, and ultimately share it with the researchers. The extracted data was then used to create the three modes of presentation tailored to each participant’s work patterns and habits. After participants willingly shared their telemetry data with us, they engaged with the three modes: viewing their data on a dashboard, reading a personalized narrative, and interacting with an AI productivity agent that had access to their telemetry data. They then assessed their experiences based on the ability to gain productivity insights or identify opportunities for productivity improvement by answering our follow-up survey. Note, the AI productivity agent primarily functioned as a ‘productivity guide,’ offering recommendations and advice to improve users’ productivity. This included suggesting strategies for time management, prioritizing tasks, and optimizing work routines. However, it is important to note that the agent did not directly perform tasks such as meeting summarization or documentation writing. While it provided guidance in these areas, the execution of tasks remained the user’s responsibility. This distinction is crucial for understanding the agent’s role in our study. In total, we had N=40 participants complete the second phase of our study. To diversify our sample, we reached out to individuals who had not participated in the first phase of the study. The average time for study completion was 30.15 minutes, and participants received a \$25 gift card as compensation. Our study exclusively makes use of OpenAI’s  (OpenAI, 2023b) GPT-4 API.
@@
Telemetry Data Collection & Challenges: Our scraper tool employs Selenium (Selenium, 2023), a testing library, for browser automation to navigate and capture data from Viva Insights pages. The system’s reliance on browser automation introduced several complexities: the scraping tool’s functionality depended on OS compatibility and browser versions, requiring careful packaging of appropriate drivers (64-bit vs 32-bit) for different user environments. We encountered occasional timeout issues during processing, necessitating robust error handling. Since these pages are dynamically generated, we opted for a more error-resilient approach by utilizing GPT-4 to extract information, rather than relying on traditional HTML tag-based scraping. However, this approach presented its own technical challenges, particularly regarding token limitations of the GPT-4 API. Our initial approach of passing entire HTML pages to GPT-4 frequently exceeded token limits. To address this, we implemented preprocessing steps using Python libraries to remove unnecessary HTML elements and select only relevant div tags before querying GPT-4. The streamlined source was then sent to GPT-4 with specific prompts to identify the desired information. The integration of telemetry data with GPT-4 required careful prompt engineering to ensure consistent interpretation and reference. We structured the system prompt to include telemetry data in a standardized format, with clear delineation between different metrics (meetings, emails, focus time, etc.) to enable reliable data access. Since users had varying availability of telemetry data, we had to identify and focus on metrics that were commonly available across users. Our prompts were designed to explicitly exclude unavailable data points to prevent inconsistent or erroneous responses. To address these variations, we developed a template based on JSON that was used to dynamically complete the telemetry data portion of the system prompt, ensuring all necessary context was preserved regardless of data volume. Please refer to Appendix C and Appendix D for the comprehensive list of system prompts and user prompts we provided to GPT-4 for extracting and transforming the desired information. The telemetry information obtained from Viva Insights can be found in Table 1. It is important to note that Viva Insights calculates these insights based on the data from the prior month. Despite these various technical challenges, we successfully implemented web systems for both the GPT-4 agent and dashboard that provided seamless user experiences without visible errors, though with some latency in response times.
@@
Designing Productivity Solutions: Having addressed these technical challenges, we proceeded with designing three distinct productivity solutions: an AI productivity agent (also referred to as an interactive chat agent), a personalized narrative (generated via GPT-4), and a traditional visual dashboard. We began by integrating telemetry data directly into the AI agent’s system prompt, a crucial step that empowered the agent with the ability to analyze and respond based on individual user behaviors and patterns. This data-driven approach ensured that each participant received feedback and suggestions that reflected their unique work habits. We then drew from the insights gathered during the first phase of our study. The preferences and expectations participants shared during this phase became the foundation for designing the system prompt of the AI agent. Parallel to the development of the AI agent, we embarked on creating an AI-based personalized narrative, an alternative means of communicating productivity insights. This narrative was also crafted to align with the specific preferences participants had voiced. Beyond just these preferences, we introduced a structured format to this narrative, ensuring as consistent delivery of insights as we could across participants. The narrative format presented information in a structured, story-like manner, providing users with an overview of their productivity patterns and suggesting improvements. In contrast, the interactive agent engaged users in a conversational interface, offering personalized guidance based on user inputs and queries. This interactive nature allowed for a more tailored and dynamic exchange, though it was limited to offering advice rather than executing tasks directly. To perfect the prompts for both the AI productivity agent and the personalized narrative, we engaged in multiple cycles of testing and refinement. Each iteration was an opportunity to align more closely with user needs and preferences. Refer to Appendix  E for the full prompts used for the productivity agent and personalized narrative. For the development of the dashboard visualization, we recognized that many participants had established familiarity with Viva Insights, so we aimed to make our dashboard’s visual presentations echo those of Viva Insights. While GPT-4 was used to extract the telemetry data that populated the dashboard, the dashboard itself was a traditional visualization tool using standard charts and metrics rather than involving any generative AI components. We designed a range of visual representations, from bar charts that provided at-a-glance insights to pie charts that broke down data into comprehensible segments. Furthermore, we chose to present certain telemetry data numerically, ensuring users had access to precise figures. Figure 3 presents the AI productivity agent along with the other two presentation modes.
@@ 4.1. Exploring User Preferences for Productivity Agents
Most of the participants, 169 (60.4%), expressed their preference for the agent to provide proactive suggestions to enhance productivity. In contrast, 77 participants (27.5%) wanted the agent to provide assistance and suggestions only when explicitly asked for help. The agent’s preferred personality traits are depicted in Figure 5(a); the most common preferences were: knowledgeable (88.9%), trustworthy (75.0%), and transparent (68.2%). Similarly, Figure 5(b) represents the social role that people desired for the agent. The top choices were that of an assistant (76.4%) and expert (56.1%). The supervisor role was the least preferred, with only 4.6% of participants opting for it.
@@
A significant majority of participants, 209 (74.6%), expressed that they would prefer a moderate to high level of personalization in the productivity agent. This enables the agent to provide highly personalized advice tailored to their specific needs, considering various aspects of their work habits and preferences. In terms of privacy, 152 participants (54.3%) felt either somewhat comfortable or completely comfortable sharing their personal information with the productivity agent. However, 58 participants (20.7%) expressed discomfort in sharing their personal information. Interestingly, a majority of participants, 205 (73.2%), were either somewhat comfortable or completely comfortable sharing high-level telemetry data for enhancing the quality and guidance provided by the productivity agent.
@@
Productivity Perspectives Participants felt productive when making progress on scheduled tasks, solving unexpected tasks, and helping colleagues Agent designed to offer strategies for task management, prioritizing, and collaborative assistance. It is developed to help users achieve an optimal work-life balance and improve their productivity.
@@
Communication Style Preference for a tone that balances casual and formal styles; adaptability in tone based on task and situation; direct and straightforward communication rather than chatty Agent uses a professional yet friendly tone, adaptable to the context and user’s current task. It is designed to communicate concisely without being overly chatty or verbose.
@@
Approach Desire for proactive suggestions and assistance to enhance productivity Agent designed to provide proactive productivity suggestions without being too intrusive and pushy
@@
Personality Traits Valued traits included being knowledgeable, trustworthy, and transparent. Agent developed to demonstrate these traits consistently across interactions.
@@
Personalization Preference for a moderate to high level of personalization, balancing with privacy concern Agent uses high-level telemetry data to personalize suggestions while ensuring user privacy
@@
Negative Traits to Avoid Avoidance of traits like being nagging, aggressive, pushy, overly casual, judgmental, or patronizing. Agent prompted to avoid these negative traits, focusing on supportive and informative interactions.
@@
4.2. Design of the AI Productivity Agent Informed by User Preferences and Telemetry Data
The development of our GPT-4-powered productivity agent focused on two key aspects: effective integration of telemetry data for personalization and alignment with user preferences identified in Phase 1 (Section 4.1). The agent’s system prompt was dynamically constructed to include the user’s telemetry data, structuring it into clear sections about meeting habits, communication patterns, focus time, and work-life balance metrics. We formatted the telemetry data to enable the agent to make specific comparisons and identify patterns, such as “Your meetings are 45% longer than the average” or “You have 11 days with after-hours work in the past month.” The agent was designed to reference this data explicitly in its responses, grounding its recommendations in the user’s actual work patterns.
Based on Phase 1 findings, we configured the agent to maintain a professional yet friendly tone (preferred by 35.4% of survey participants), provide proactive suggestions while respecting user autonomy (aligned with 60.4% of participants’ preferences), and demonstrate traits of being knowledgeable and trustworthy (preferred by 88.9% and 75.0% of participants respectively). A summary of these findings and their corresponding implementations in the AI agent’s design is presented in Table 2. The table illustrates how each identified theme from the survey was translated into specific features of the AI productivity agent. The agent was specifically instructed to avoid characteristics identified as negative in our survey, such as being pushy or overly casual (See agent’s system prompt in Appendix E).
Additionally, personalization emerged as a crucial aspect, with users expressing a need for tailored recommendations balanced against privacy concerns. Accordingly, the agent leverages high-level telemetry data to offer personalized suggestions, carefully designed to respect user privacy. When a user asks about improving their communication habits to boost productivity, the agent analyzes their personal communication metrics from the telemetry data, identifies specific patterns, and provides recommendations grounded in their actual behavior. For example, as shown in Appendix I Figure 9(b), when a user asks, “What can I improve in my communication habits to boost productivity?”, the system provides personalized recommendations based on their specific work patterns. The agent identifies that the user engages heavily in email and communication during their peak productivity hours (12-3 PM) and suggests redistributing these activities to less productive times. The system also recognizes work-life balance issues by noting after-hours communication occurring in 11 out of 28 days. This example demonstrates how telemetry data enables meaningful personalization through analysis of broader work patterns—identifying meeting scheduling trends, communication habits, focus time preferences, and work-life balance patterns—to provide targeted recommendations for productivity improvement.
@@ 4.2. Design of the AI Productivity Agent Informed by User Preferences and Telemetry Data
Figure 6. Overall Experience: Figure (a) illustrates the participants’ overall experience ratings for three data presentation methods: Personalized Narrative, AI Productivity Agent, and Dashboard. Figure (b) compares the ratings of these methods across various criteria. We performed the Kruskal Wallis test to analyze the differences between the three groups in Figure (b). Criteria without any statistically significant differences (i.e., p-value > 0.10) are marked with a crosshatch shading pattern - a series of intersecting lines drawn at an angle to create texture.
@@
Participants spent an average of 60.58 seconds (with a standard deviation of 38.98 seconds) interacting with the dashboard. In comparison, they spent considerably more time on the AI productivity agent, with an average duration of 315.48 seconds (standard deviation: 272.23 seconds). The personalized narrative had a similar interaction time to the dashboard, with an average of 65.21 seconds (standard deviation: 55.39 seconds). After interacting with each method, participants were asked to rate their overall experience on a six-point Likert scale ranging from Very Poor to Excellent. As shown in the stacked bar graph in Figure 6(a), 26.2% of participants rated the Dashboard as excellent, while only 11.9% and 9.5% rated Excellent for the productivity agent and personalized narrative, respectively. Combining the top three experiences (excellent, very good, and good), 81% considered the Dashboard good or above, 66.7% considered the AI productivity agent good or above, and 64.2% considered the personalized narrative to be good or above. To compare user ratings across the three methods, we conducted a chi-square test of independence. Given our relatively small sample size, we used a significance level of $\alpha$ = 0.10 to account for reduced statistical power. The analysis revealed statistically significant differences in the distribution of ‘Excellent’ ratings across conditions ($\chi$2 = 5.16, p-value = 0.07), suggesting that participants had different preferences when giving ‘Excellent’ ratings to the three conditions. We did not find any significant differences in the distribution of other rating categories (Very Good, Good, Fair, Poor, Very Poor) across the methods. Following the broader question, participants were asked to rate their experience on specific criteria for each method using a five-point Likert scale from Strongly Disagree (1) to Strongly Agree (5). In Figure 6(b), a heatmap displays the average scores obtained by each method for these questions. Most darker blocks (in blue) are centered on the Dashboard, indicating higher average scores for the ease of exploring and understanding data, the use of the presentation method in real-life situations, and enjoyment in interaction. The boxes with crosshatch shading patterns (intersecting lines) indicate that the differences were not statistically significant (Kruskal Wallis test; p-value > 0.10) among the three methods for that particular criterion. The productivity agent scored better in only one area: engagement during data exploration. The personalized narrative performed the worst with the lowest scores among the three. Regarding anticipated usage frequency, we show in Figure 7(a) that 42% of participants selected weekly for the dashboard and 30% selected weekly for the productivity agent, which was the majority score for both groups. For personalized narrative, 28% of participants said that they would use it weekly and the same amount also said they would rather use it monthly. 30% of the participants said they would use the personalized narrative occasionally. However, we did not obtain any statistically significant differences in the distribution of usage frequencies across the three methods. Lastly, participants ranked the three methods in order of preference (1 being the most preferred and their first choice, 3 being the least preferred and their last choice). As shown in Figure 7(b), 55% rated the Dashboard as their most preferred method, followed by 27.5% preferring the productivity agent and 17.5% preferring the personalized narrative. Over half of the participants (52.5%) rated the personalized narrative as their least preferred method. A chi-square test of independence revealed significant differences in order of preference across the three methods for rank 1 i.e. most preferred method ($\chi$2 = 13.57, p-value < 0.01). The least preferred method i.e. rank 3 also showed significant differences ($\chi$2 = 14.47, p-value < 0.01), but not rank 2.